The shareholders of Bread Company have voted in favor of a buyout offer from Butter Corporation. Information about each firm is given here: Bread Butter Price-earnings ratio 16 33 Shares outstanding 96,000 330,000 Earnings $ 190,000 $ 950,000 Bread's shareholders will receive one share of Butter stock for every three shares they hold in Bread. a-1. What will the EPS of Butter be after the merger? (Do not round intermediate calculations and round your answer to 3 decimal places, e.g., 32.161.) a-2. What will the PE ratio be if the NPV of the acquisition is zero? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.) b. What must Butter feel is the value of the synergy between these two firms?
Answer #1

a-1)

Here bread company will get merged into butter company

Bread's shareholders will receive one share of Butter stock for every three shares they hold in Bread

hence number of shares to be issued by butter company = 96000/3 = 32,000 shares

Hence after merger shares of butter company = 330,000 + 320,00 = 362000 shares

After merger earnings of combined entity = Bread's Earning before merger + Butter's earning before merger

= 190,000+950,000

=11,40,000 $

Thus EPS of butter after merger = After merger earnings/ After merger number of shares

= 1140,000/362000

=3.149 $

a-2) When there is 0 NPV acquisition , there will be no effect on share price

Thus share price of butter = PE ratio x EPS

= 33 x (950000/330000)

= 33 x 2.879

= 95$

PE ratio after merger = Price per share of butter company / EPS of butter company after merger

=95/3.149

= 30.17 times

b) Value of bread to butter must be market value of bread since NPV is 0

Market value of bread = PE ratio X Earnings

= 16 x 190,000

= 3040,000$

Cost of acquisition = No of shares offered x market price per share of butter

= 32000 x 95

= 3040,000$

Thus NPV of acquisition is 0

Although there is no economic value of takeover, it is possible that butter is motivated to purchase bread for reasons other than financial reasons like

Reduce in competition

Increase in market share, etc